OpenStudy (anonymous):

Find the value of x. Round to the nearest tenth. Answer A. 62 B. 25.5 C. 27.8 D. 25

OpenStudy (anonymous):



OpenStudy (phi):

do you know pythagoras? a^2 + b^2 = c^2 where c is the length of the longest side (opposite the right angle)

OpenStudy (anonymous):

use sin

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it 27

OpenStudy (anonymous):

sin-1x=7/15

OpenStudy (anonymous):

yes its 27.81813928